OpenStudy (anonymous):

f(a)=(7/(a-6))-(84/(a^2-36)) evaluate the limit where a approaches 6 f(a)=? please explain the steps

OpenStudy (anonymous):

lim(7(a+6)-84)/(a^2-36)=lim(7a-42)/(a^2-36)=lim7/(a+6)=7/12, ok ?

OpenStudy (anonymous):

I don't think that's right

OpenStudy (anonymous):

0/0 ?

OpenStudy (anonymous):

limit a -> 6

OpenStudy (anonymous):

are u mathtype and word 2003 ?

OpenStudy (anonymous):

lim (7/(a-6))-(84/(a^2-36)) =lim((7/(a-6))-(84/((a-6)(a+6))) =lim ((7(a+6)-84)/((a-6)(a+6))) =lim ((7a+42-84)/((a-6)(a+6))) =lim ((7a-42)/((a-6)(a+6))) =lim(7(a-6)/((a-6)(a+6))) =lim (7/(a+6)) =7/12

OpenStudy (anonymous):

ok